### v2

# HebID: Detecting Social Identities in Hebrew-language Political Text

Abstract
Political language is deeply intertwined with social identities. While social identities are often shaped by specific cultural contexts, existing NLP datasets are predominantly English-centric and focus on coarse-grained identity categories. We introduce HebID, the first multilabel Hebrew corpus for social identity detection. The corpus contains 5,536 sentences from Israeli politicians’ Facebook posts (Dec 2018–Apr 2021), with each sentence manually annotated for twelve nuanced social identities (e.g., Rightist, Ultra-Orthodox, Socially-oriented) selected based on their salience in national survey data. We benchmark multilabel and single-label encoders alongside 2B–9B-parameter decoder LLMs, finding that Hebrew-tuned LLMs provide the best results (macro-$F_{1}$ = 0.74). We apply our classifier to politicians’ Facebook posts and parliamentary speeches, evaluating differences in popularity, temporal trends, clustering patterns, and gender-related variations in identity expression. We utilize identity choices from a national public survey, comparing the identities portrayed in elite discourse with those prioritized by the public. HebID provides a comprehensive foundation for studying social identities in Hebrew and can serve as a model for similar research in other non-English political contexts. [^1] https://github.com/guymorlan/hebid/
HebID: Detecting Social Identities in Hebrew-language Political Text
Guy Mor-Lan1, Naama Rivlin-Angert1, Yael R. Kaplan2, Tamir Sheafer1, Shaul R. Shenhav1 1 The Hebrew University of Jerusalem, 2 The Open University of Israel guy.mor@mail.huji.ac.il
1 Introduction
Social identities—such as political ideology, religious affiliation, or demographic group membership—are powerful drivers of political behavior and public discourse (Tajfel and Turner, 1979). Yet existing NLP resources for identity detection remain almost entirely English‐focused, single‐label, and rely on coarse-grained categories (e.g., party or ethnicity). In this work, we introduce HebID, the first publicly released Hebrew dataset for fine‐grained, multilabel social identity detection in political text, grounded in both domain expertise and large‐scale survey evidence.
We frame the task as a sentence-level classification problem: the goal is to determine which, if any, of twelve social identities are being positively expressed or invoked within a given sentence.
The choice of social identities is empirically grounded in survey data. We utilized 12 waves of a national survey (N = 1,769) each of which included questions designed to identify the social identities most salient to Israeli citizens (e.g., Rightist, Ultra-Orthodox, Socially-oriented). These survey‐derived categories were then used to guide the annotation of 5,536 sentences sampled from Facebook posts by Israeli politicians (December 2018–April 2021), ensuring that our labels reflect real‐world identity salience.
We benchmark a suite of modern architectures on the Facebook data—(i) multilabel encoder models; (ii) single‐label encoder models; and (iii) 2B–9B‐parameter decoder LLMs, finding that Hebrew‐tuned decoder models (specifically, DictaLM2.0) achieve the highest macro‐F1 (0.743), outperforming encoder‐only baselines by over six points. We then assess cross‐genre generalization by applying our best model to 500 parliamentary speech excerpts, achieving a comparable macro‐F1 of 0.72.
Finally, we link three complementary sources in our analysis of how identities behave: politicians’ Facebook posts, parliamentary speeches from the Israeli parliament (Knesset), and survey responses that reflect public identification. We (1) compare identity prevalence and correlations between social media, parliamentary speech, and the public; (2) document how identity discourse surges around election cycles; (3) uncover coherent “bundles” of co‐occurring identities; and (4) quantify gender‐related variation in identity expression.
Our contributions are:
  • Dataset and survey-grounded methodology: We introduce a framework for creating identity corpora by linking survey data to text annotation, and release HebID, the resulting Hebrew dataset.
  • Comprehensive benchmarks: We evaluate a range of encoder-only and decoder LLMs, establishing strong baselines for the task.
  • Cross-genre evaluation: We demonstrate our best model’s generalization to parliamentary speech, confirming its robustness.
  • External validation: We validate our classifier against an external expert survey (CHES-Israel), showing its alignment with party policy positions.
  • Sociolinguistic analysis: We use the classifier to reveal novel insights into identity dynamics across social media, parliamentary speeches, and public surveys.
Taken together, these contributions offer value on three fronts. For computational social scientists, HebID provides a novel tool for analyzing political discourse. For the NLP community, our work establishes a challenging new benchmark for multilabel classification in a low-resource language. Finally, our survey-grounded annotation framework serves as a methodological template for developing similar culturally-aware resources in other non-English contexts.
2 Related Work
Automatic analysis of social identity language in political text intersects several NLP subfields: group reference detection, framing and stance classification, and ideological position inference. However, existing resources remain limited in granularity, language coverage, and multilabel capacity, leaving a clear gap that our Hebrew dataset fills.
Group reference detection.
Early work extended named‐entity recognition to capture social groups as entities. Zanotto et al. (2024) introduce GRIT, annotating Italian news and parliamentary text for spans referring to demographic, national, or partisan groups, and fine‐tune BERT to identify them. Licht and Sczepanski (2024) apply a similar span‐labeling approach to British parliamentary debates, quantifying how often politicians mention particular social groups. These studies demonstrate the feasibility of automatic group mention detection, but remain single-label and limited to explicit group mentions.
Framing and stance.
Beyond mentions, understanding how groups are portrayed is crucial. The Us vs. Them corpus (Huguet Cabot et al., 2021) annotates Reddit comments for target group, stance (supportive vs. hostile), and emotion, using a multi‐task RoBERTa model. Card et al. (2022) study 140 years of U.S. congressional speeches on immigration, combining sentiment classification with custom frame lexica to reveal evolving and polarized frames. The Media Frames Corpus (Card et al., 2015) provides frame annotations (including cultural identity frames) for news articles. These resources focus on single‐label stance or framing, and predominantly English texts.
Ideological position inference.
Separate but related is the task of inferring latent political ideology. Thomas et al. (2006) use SVMs on floor debate transcripts to classify support/opposition. Iyyer et al. (2014) develop neural networks to predict left/right alignment of speeches. More recent prompt‐based methods with LLMs (Mishra et al., 2023) achieve zero‐shot ideological scoring. While these approaches infer broad ideological leanings, they do not detect explicit identity mentions or support multilabel identity categories.
In‐group/out‐group rhetoric.
Discourse‐level signals such as pronoun clusivity and coded language reveal populist identity appeals. Rehbein and Ruppenhofer (2022) annotate inclusive vs. exclusive uses of “we” in German parliamentary debates and train classifiers to disambiguate referents, highlighting us‐versus‐them framing. Mendelsohn et al. (2023) curate a dogwhistle lexicon and show that pretrained models struggle to detect covert slurs without knowledge grounding. These works emphasize the need for high‐precision, context‐aware annotation, but do not cover multilabel identity categories across many classes.
Gaps and Our Contributions.
Current identity‐focused corpora are typically single‐label, English‐centric, and restricted to broad political categories or explicitly stated group mentions. They rarely offer the fine‐grained, multilabel annotations needed to capture the complexity of real‐world political speech, and they do not cover non‐Latin scripts or languages such as Hebrew. We introduce the first publicly available Hebrew dataset for social identity detection, comprising 5,536 sentences from politicians’ Facebook posts annotated with twelve distinct identities—each grounded in survey‐measured salience and expert‐defined categories. By providing multilabel annotations across a rich set of ideologically, religiously, and socio‐economically relevant identities, our resource enables more nuanced analyses of how political actors deploy identity language than has previously been possible in any non‐English setting.
3 Annotating Social Identities
3.1 Panel Survey
To select social identities for annotation, we utilized a combination of experts and survey instruments. We utilized a representative 12-wave panel survey of the Jewish population in Israel (N = 1,769), conducted between January 2019 and April 2021 (Dvir-Gvirsman et al. 2022; see also Rivlin-Angert et al. 2025). [^2] Similar to other panel surveys in which local sample vendors were unable to include re-interview samples of Palestinian citizens in sufficient numbers (e.g., Gidron et al. (2022)), the panel survey did not include this population, reflecting a known deficiency in Israel’s survey sampling market. This time period contains four Israeli elections held in quick succession, providing a unique opportunity to investigate political dynamics in a condensed time frame. In 12 survey waves, respondents were asked to select up to three identities they identify with, out of a list of 28. These 28 identities were chosen by a panel of experts in Israeli politics as reflecting a broad spectrum of prevalent social identities, spanning ideological, value-based, religious, national, socio-demographic, and economic dimensions (e.g., Conservative, Leftist, Nationalist, LGBTQ+. See appendix A for a full list).
For inclusion in the textual annotation, we selected the 12 identities that emerged as most salient in the panel survey, each consistently surpassing a 5% selection threshold in the first five waves. These identities are Rightist, Leftist, Conservative, Liberal, Socially-oriented, Capitalist, Zionist, Palestinian, Honest, Security-oriented, Ultra-orthodox, and Democrat. [^3] The 5% criterion was jointly applied to identities that respondents identify with and identities they disapproved of in a separate survey item. For more information on the survey methodology, see appendix B.
3.2 Annotation Scheme
Based on this selection, we developed an annotation scheme for the expression of social identities in text. Identities were annotated only when referenced in a positive manner; negative or oppositional mentions were excluded. Each sentence was annotated using a multilabel scheme across the twelve selected identities.
Below is a summary of the identity definitions used in the annotation scheme, accompanied by examples from the dataset translated from Hebrew (full definitions appear in appendix C). Some examples were abridged for brevity.
Liberal:
[…]
Conservative:
[…]
Democrat:
[…]
Leftist:
[…]
Rightist:
[…]
Capitalist:
[…]
Socially-oriented:
[…]
Zionist:
[…]
Security-oriented:
[…]
Honest:
[…]
Palestinians and Arab Citizens of Israel:
[…]
Ultra-orthodox:
References to the Jewish ultra-Orthodox lifestyle, including the education system, gender segregation, and exemption from military service, as well as mentions of ultra-Orthodox political parties and leadership.
[…]
3.3 Co-occurring Identities
Our annotation scheme is multi-label, allowing a single sentence to express several social identities simultaneously. Each annotated sentence is evaluated with respect to all included identities. For example, the following sentence expresses Democrat, Leftist, and Honest social identities:
[…]
Similarly, this sentence expresses both a Rightist and a Liberal identity:
[…]
4 Dataset
To sample sentences for annotation, we compiled a corpus of Facebook posts by Israeli members of parliament, political parties, and viable party candidates, posted during the same timeframe as the survey (Dec. 2018 to Apr. 2021). This corpus comprises 64K posts containing 375K sentences. Our annotated dataset contains 5,536 Hebrew-language sentences sampled from 4.8K unique Facebook posts. Each sentence is annotated with 12 binary annotations according to whether it expresses a given identity. The dataset is split into a training set (70%), validation set (15%) and test set (15%).
4.1 Inter-coder Reliability
The sentences were annotated by two of the authors. We evaluate inter-coder reliability on a sample of 304 sentences. The results, given in Table 1, indicate a mean Cohen’s $\kappa$ of 0.77. Disagreements were resolved via consultation with all authors.
Category Percent Agreement Cohen’s $\kappa$
Conservative 96.7% 0.705
Rightist 94.4% 0.788
Democrat 91.4% 0.703
Honest 96.7% 0.782
Capitalist 98.4% 0.792
Ultra-Orthodox 98.7% 0.708
Socially-oriented 96.1% 0.841
Liberal 96.1% 0.841
Leftist 94.1% 0.761
Security-oriented 96.4% 0.736
Palestinian 98.4% 0.792
Zionist 96.1% 0.778
Mean 96.1% 0.769
Table 1: Inter‐Annotator Agreement (IAA) scores
4.2 Descriptive statistics
The number of positive labels per identity ranges between 129 (Ultra-Orthodox) and 703 (Rightist), with a mean of 413 (Table 2). 37.5% of sentences express no identity, 41.1% express one identity and 21.4% express two or more identities (Table 3).
Category Count
Rightist 703
Liberal 629
Socially-oriented 567
Democrat 562
Leftist 546
Zionist 368
Honest 357
Security-oriented 346
Conservative 297
Capitalist 230
Palestinian 225
Ultra-Orthodox 129
Mean 413.25
Table 2: Number of Positive Instances
# Positives Row Count Percent
0 2,078 37.54%
1 2,275 41.09%
2 919 16.6%
3 218 3.94%
4 39 0.7%
5 6 0.11%
6 1 0.02%
Table 3: Number of Positive Labels per Sentence
5 Modeling and Experiments
We model the task of identifying social identities from text using four types of training setups:
First, we establish Baseline models by training separate Logistic Regression and LinearSVC classifiers for each label using TF-IDF features. Then, Multilabel encoder-models are fine-tuned to jointly learn the 12 identity labels. Single-label encoder models are fine-tuned separately for each label. For this training, we utilize the best performing base model in the multilabel encoder setup. Finally, Decoder LLMs are fine-tuned to generate a comma-separated list of all applicable Hebrew language labels for the given sentence.
We examine a set of encoder models and LLM decoders in the 2B–9B parameter range. For encoders, we use the multilingual mBERT (Devlin et al., 2019), and Hebrew-targeted encoders AlephBERT (Seker et al., 2022), HeRO (Hebron et al., 2023) and the base and large variants of DictaBERT (Shmidman et al., 2023). All encoder models are trained for up to 10 epochs with three learning rates (1e-5, 3e-5, 5e-5), and three types of loss (default, positive weight and focal loss), choosing the best performing checkpoint on the validation set.
For decoder LLMs, we use Gemma 2 in the 2B and 9B variants (Gemma Team, 2024), Qwen3-8B (Qwen Team, 2025), and the Hebrew-targeted DictaLM2.0 (Shmidman et al., 2024). All decoders are fine-tuned with QLORA (Dettmers et al., 2023) for up to 5 epochs with two learning rates (1e-5, 1e-4).
Macro-averaged metrics Per-label F1 scores
Model P R F1
Security-
oriented
Capitalist
Conserva
-tive
Democrat
Ultra-
Orthodox
Socially-
oriented
Liberal
Palestinian
Leftist
Rightist
Honest
Zionist
Decoder‐only
DictaLM2.0 0.740 0.751 0.743 0.705 0.805 0.675 0.754 0.653 0.723 0.724 0.852 0.750 0.765 0.816 0.700
Gemma-2‐9B 0.717 0.698 0.705 0.645 0.759 0.684 0.753 0.591 0.662 0.749 0.758 0.671 0.737 0.804 0.650
Gemma-2‐2B 0.620 0.631 0.624 0.560 0.723 0.575 0.757 0.385 0.634 0.673 0.716 0.609 0.615 0.680 0.557
Qwen‐8B 0.665 0.463 0.542 0.605 0.508 0.308 0.700 0.410 0.541 0.568 0.625 0.516 0.515 0.650 0.554
Multilabel encoders
DictaBERTLarge 0.677 0.680 0.678 0.660 0.667 0.561 0.750 0.667 0.688 0.689 0.831 0.588 0.670 0.716 0.645
[…]
AlephBERT 0.628 0.692 0.657 0.673 0.675 0.526 0.738 0.604 0.627 0.664 0.783 0.620 0.641 0.718 0.618
HeRo 0.608 0.693 0.647 0.667 0.575 0.587 0.730 0.528 0.648 0.642 0.776 0.659 0.638 0.720 0.589
mBERT 0.573 0.553 0.552 0.483 0.467 0.527 0.708 0.356 0.568 0.635 0.632 0.531 0.502 0.653 0.566
Single‐label encoders
DictaBERTLarge 0.643 0.689 0.659 0.672 0.659 0.529 0.731 0.508 0.708 0.694 0.783 0.615 0.636 0.721 0.652
Baselines
LinearSVC 0.398 0.339 0.361 0.400 0.170 0.300 0.490 0.230 0.430 0.330 0.520 0.390 0.390 0.350 0.350
LogisticRegression 0.469 0.268 0.334 0.350 0.000 0.280 0.500 0.280 0.410 0.340 0.470 0.380 0.350 0.360 0.290
Table 4: Macro‐averaged precision (P), recall (R), F1, and per‐label F1 for all models.
Table 4 reports test set results for trained models. We observe that the most performant are decoder models, achieving the highest per-label F1 scores on all but one label. DictaLM2.0, specifically, achieves the best macro P/R/F1 scores, and best per-label scores on 8/12 identities. The performance of this model, continuously pre-trained from Mistral 7B on 100B Hebrew tokens, underscores the dependence identity detection on language specific cultural and political context.
[…]
5.1 Generalization to Parliamentary Speech
We examine whether the resulting classifier generalizes well to the parliamentary speech in the Knesset by utilizing the IsraParlTweet dataset (Mor-Lan et al., 2024). We subset speeches from the equivalent time period of the Facebook data, classify all sentences with the best performing DictaLM2.0 model, and sample 500 of the sentences for human annotation. We perform a precision-oriented test by oversampling positive predictions for each identity. The results show a macro $F_{1}$ score of 0.72, on par with the Facebook test set, indicating the model’s ability to generalize to Knesset data (for full results, see Table 5).
5.2 External Validity
We examine the external validity of the classifier by correlating social identity discourse with external party policy rankings from the CHES-Israel expert survey (Zur and Bakker, 2025), which assigns parties scores on several policy domains. For each of the 16 political parties covered by both the expert survey and our Facebook corpus in year 2021, we calculate identity discourse scores, defined as the share of sentences posted by the party’s politicians which are classified as expressing said identity.
We match closely related policy issues from the expert survey to each social identity. For example, policy issue Economic Stance is linked with both Capitalist and Socially-oriented social identities. For better comparability, in some cases we additionally construct social identity dimensions by subtracting party means for two polar identities (e.g. $\text{Rightist}-\text{Leftist}$, $\text{Capitalist}-\text{Socially-Oriented}$). We then compute correlations between the CHES party policy scores and the predicted social identity discourse scores, for the 16 political parties.
Our analysis reveals a strong alignment with the expert-coded CHES data. Out of 21 correlations, 16 are statistically significant at $p\leq 0.1$ and 13 are significant at $p\leq 0.05$. Of those 13, all fall within a strong effect size range, with absolute correlation values from $|r|=0.71$ to $0.94$. For full results, see Table 7 in the appendix.
6 Results
We utilize the three identity data sources with overlapping time frames: our corpus of Israeli politician Facebook posts; the subset of parliamentary speeches from (Mor-Lan et al., 2024) in the same time period; and the panel survey capturing the public’s identity choices. The first two sources are classified using the best performing DictaLM2.0 model. Utilizing these three data sources allows for a unique exploration of the dynamics of social identities on different platforms of elite discourse (Facebook and Knesset speeches) and the public (panel survey). We examine the following aspects: the popularity of identities; temporal trends; the correlation and bundling together of identities; and gender differences in identities.
Popularity.
In Figure 1, we compare the normalized share of each identity. The identities Socially-oriented, Rightist and Democrat are generally popular across the data sources. Interestingly, the Leftist identity is relatively less popular. While many identities receive similar proportions among the three sources, several exceptions stand out. Identities Honest and Zionist are significantly more popular among the public, whereas Socially-oriented is significantly more popular in parliament.
Examining correlations between the ranks of identities in each data source, we see that while the popularity ordering of identities is strongly correlated between the Facebook and Knesset data ($\rho$=0.87), survey identity rankings are moderately correlated with Facebook ($\rho$=0.48) and Knesset ($\rho$=0.46) data. For all ranks, see Figure  10.
Figure 1: Normalized share of identities
Temporal trends.
We examine whether identity-related discourse tends to increase before elections. Facebook posts are the most suitable data source for this analysis, as the survey panel includes a limited number of waves and Knesset meetings are in recess near election periods. We plotted the average number of identity mentions per sentence per week in the Facebook data (see Figure 2). The results show considerable fluctuation, ranging from 0.38 to 0.87 identity mentions per sentence. Notably, identity-related discourse peaks around three of the four election dates, highlighting the salience of identities during electoral competition.
Figure 3 plots the share of six identities over time for both the Facebook and survey samples (see Figure 11 for all 12). We see that the identities that peak during or near election dates are Rightist, Leftist and Democrat. However, as previously mentioned, Leftist holds a very small share in both Facebook and the survey, and thus drives a smaller part of the overall temporal change. A notable gap between elites and the public emerges for three identities. While the public decreasingly identifies with the Socially-oriented identity, it becomes significantly more prominent among politicians after the 3rd elections held in March 2020. On the other hand, the identities Honest and Democrat become more popular among the public after the 3rd election, but less popular in elite discourse. These findings demonstrate the potential of HebID to provide valuable insights into how identity-driven agendas shift in the lead-up to elections.
Figure 2: Temporal Trend
Figure 3: Per Identity Trends - Facebook and Survey
Identity bundles.
To examine which identities tend to co-occur, we perform a factor analysis on the mean levels of each identity per speaker/respondent (Figure 4). In the three sources, we see a primary dimension dividing identities into two broad groups, a left-wing group (Leftist, Democrat, Honest, Liberal, Palestinian) and a right-wing group (Rightist, Conservative, Zionist, Security-oriented, Capitalist, Ultra-orthodox). Other factors in the Facebook and Knesset data appear to reflect political sub-dimensions of identity pairs that are more tightly coupled, such as Conservative and Rightist, Zionist and Security-oriented, Liberal and Palestinian, and Honest and Democrat. In the survey data, factors beyond the first dimension appear less structured compared to those that arise in the Facebook and Knesset datasets.
Figure 4: Factor analysis
Accordingly, examination of the correlations between identities (after aggregating into speakers/respondents) shows that the survey exhibits a mean absolute correlation coefficient of 0.159, weaker than the Facebook sample (0.235) and Knesset data (0.215). For full correlation matrices, see appendix K.
Gender differences.
Do men and women differ in terms of social identities? We first examine gender differences in identity discourse by calculating gender differences in the total number of identities expressed. For each speaker/respondent, we calculate the mean number of identities expressed in a sentence or survey response, and then aggregate by gender. We see that in all data sources, women express more identities than men. However, the gap is largest in the Knesset data (0.07) and is not statistically significant in the survey data.
We then examine the gender difference per identity by subtracting the share of each identity among women from that of men, in all data sources. We see that some identities, such as Rightist, Security-oriented, Capitalist, and Ultra-orthodox lean towards men, whereas Socially-oriented significantly leans towards women across all platforms. While the gender gaps on Facebook and the Knesset are generally in the same direction, several of the survey gender gaps arise in different directions. Thus, Honest in the survey leans significantly towards women, whereas in the Facebook and Knesset data it leans slightly towards men.
Figure 5: Gender Difference per Identity
Figure 6: Gender Difference in Identity Mentions
7 Conclusion
In this work, we have introduced HebID, the first publicly available multilabel Hebrew corpus for fine-grained social identity detection in political text, grounded in large-scale survey evidence and expert consultation. Our dataset of 5,536 sentences from Israeli politicians’ Facebook posts is annotated for twelve empirically salient identities, ensuring both linguistic and sociological validity. We benchmarked a range of models—multilabel and single-label encoders as well as 2B–9B-parameter decoder LLMs—and demonstrated that Hebrew-tuned decoder models (e.g., DictaLM2.0) achieve the highest macro-F1 (0.743), significantly outperforming encoder-only baselines. We further validated cross-genre robustness by applying our best model to 500 Knesset speech excerpts, matching social-media performance (macro-F1 = 0.72).
By linking three complementary sources — Facebook posts, parliamentary speeches, and a national survey — we revealed systematic differences in identity prevalence, temporal surges around elections, bundles of co-occurring identities, and gendered patterns of identity expression. These findings showcase the potential of our framework to uncover nuanced political dynamics and address substantive questions in the social sciences.
The HebID annotation scheme and corpus provides a comprehensive foundation for future research on identity discourse in non-English political contexts and paves the way for comparative studies across political systems. Specifically, this work provides a new tool for social scientists studying political communication, a new benchmark for NLP researchers working on Hebrew and multilabel classification, and a methodological template for creating similar resources cross-culturally.
Limitations
While HebID represents a significant step forward in Hebrew political text analysis, several limitations should be acknowledged:
  • Temporal and platform scope: Identity discourse is dynamic and may have evolved beyond the period captured in our data. Additionally, other platforms—such as Twitter and news media—are not represented, leaving out potentially important dimensions of identity expression and change over time.
  • Survey population: The panel survey sampled only Jewish citizens of Israel. Identities and salience patterns may differ among non-Jewish citizens, including Palestinians and non-Jewish immigrant communities.
  • Annotation granularity: Although multilabel, our scheme relies on twelve categories selected via a 5% survey threshold. Less frequent but potentially important identities were excluded, and negative or critical references to an identity are not captured.
  • Model biases: Our classifiers inherit biases present in both the training data and the pretrained language models (e.g., under- or over-representation of certain groups). Performance may degrade on dialectal text, informal registers, or in hostile political discourse.
  • Cross-genre validation: The Knesset evaluation, while indicative of robustness, is based on a limited sample of 500 human-annotated sentences drawn from Knesset Plenum protocols. A broader evaluation across other legislative bodies or timeframes is needed to fully assess generalization.
  • Methodological considerations: Survey responses and text analysis offer complementary but distinct measurement approaches. Survey data provide insights into self-reported identity salience, while text analysis reveals patterns of expressed identity discourse. These methodological differences suggest caution when comparing panel survey results with computational text analysis findings.
Future work should extend HebID to additional platforms and populations, refine the annotation taxonomy to include emerging identities, and explore methods to mitigate model bias in identity detection tasks.
Acknowledgments
This work was supported by the Israel Science Foundation (Grant no. 2315/18). We thank Vered Porzycki for assistance in conceptualizing the definitions of the annotated identities; Vered Porzycki, Gal Ron, Avishai Green, Dror Markus and Noa Amir for invaluable assistance in a preliminary pilot study; and Effi Levi for his very useful and insightful advice.
Appendix A Full list of social identities in Survey
  1. 1.
    Ashkenazi (Jewish of European descent)
  2. 2.
    Capitalist
  3. 3.
    Conservative
  4. 4.
    Democrat
  5. 5.
    A man of faith
  6. 6.
    Honest
  7. 7.
    Humanist
  8. 8.
    Israeli
  9. 9.
    Jewish
  10. 10.
    Leftist
  11. 11.
    LGBTQ+
  12. 12.
    Liberal
  13. 13.
    Lower class
  14. 14.
    Man
  15. 15.
    Middle class
  16. 16.
    Mizrahi (Jewish of Middle Eastern/North African descent)
  17. 17.
    Nationalist
  18. 18.
    Palestinian
  19. 19.
    Religious
  20. 20.
    Rightist
  21. 21.
    Secular
  22. 22.
    Security-oriented
  23. 23.
    Socialist
  24. 24.
    Socially-oriented
  25. 25.
    Ultra-Orthodox (Haredi)
  26. 26.
    Upper class
  27. 27.
    Woman
  28. 28.
    Zionist
Appendix B Survey Methodology
The survey we utilize is an online panel survey conducted between January 2019 and April 2021, using a representative sample of the adult Jewish population in Israel. Participants were recruited through iPanel, the largest online survey firm in Israel, and received gift cards as compensation. To ensure representativeness, respondents were selected using demographic quotas based on key population parameters: age, gender, education, geographic location, and religiosity. A total of 1,769 individuals completed the first wave in January 2019, and 878 panelists completed the relevant waves for this study.
To measure identity salience, respondents were asked the following question: "People feel a sense of belonging to some groups in society and a sense of distance from others. Below is a list of groups. Please select up to three groups that best define who you are, groups that make you feel a sense of belonging, identification, and pride, and that you would like to see properly represented in the Knesset."
The possible selection options are the 28 identities appearing in section A of the appendix, a list developed by experts in Israeli politics to ensure relevance and comprehensiveness.
In total, the panel included 14 waves and an additional minor wave in which only several question items were repeated. We utilize 12 of the waves in which the identity salience question appears (11 major waves and the minor wave).
Appendix C Identity Definitions: Codebook
Dear Coder,
Below are twelve identity definitions for classification. Please follow the following guidelines:
  1. 1.
    Classify Based on Definitions. Assign identity categories only to statements that align with the definitions provided.
  2. 2.
    Multiple Classifications. If a statement may fit more than one identity category. Please assign it to all relevant identity categories accordingly.
  3. 3.
    Positive Associations Only. Please note that the identity categories reflect a positive association. Do not classify statements that oppose an identity (e.g., criticism of capitalism, liberalism, Ultraorthodox) under that category.
  4. 4.
    The Speaker Identity Is Irrelevant. Classification is based on the content of the statement, not who says it (e.g., a non-Palestinian can make a statement classified under “Palestinians and Arab Citizens of Israel” if they present a Palestinian perspective).
  5. 5.
    Not all statements will be classified. Some statements may not fit any of the twelve identity categories mentioned in the codebook. In such cases, do not force a classification.
Identity definitions:
Capitalist. Support for private enterprises, free markets and economics (including the stock market), wealth accumulation, trade-related topics, deregulation policies, tax reductions, limiting government involvement in the economy or social services. General references to economic growth, market openness, or business encouragement—without mention of state intervention—will be labeled under this identity.
Socially-oriented (Hevrati). Support for social justice and welfare-oriented policies, including reference to poverty in Israel, job security, healthcare, government-funded education, social organizations, aid initiatives (e.g., food donations), social policy (e.g., housing assistance, support for disadvantaged populations), social security benefits, infrastructure investment, prioritizing local business over international ones, and environmental policy.
Conservative. Endorsement of opposition to change, support for the integration of religion and state, preservation of traditional values, promotion of anti-liberal values, and advocating for the reduction of judicial authority. Note: This identity includes reference that endorse Jewish traditions and opposition to anti-religious coercion.
Liberal. Advocacy for equality, human and civil rights, pluralism, separation of religion and state, freedom of religion, the promotion of universal values, protection of minority or LGBTQ+ rights, and support for the judicial system.
Note that:
  1. 1.
    References relating to inequality in the burden of civic duties will be labeled under this identity.
  2. 2.
    This identity does not include references to the rule of law.
Democrat. Emphasis on democratic values and procedures, such as fair elections, the rule of law, institutional checks and balances (“rules of the game”), and the defense of democracy as a political principle. Includes explicit references to democracy or portraying Israel as a democratic state.
Honest (Yeshar Derech). References related to corruption and investigations involving public officials, honesty, integrity, and ethical conduct in public service. Includes mentions of improper immunity, criticism of self-serving behavior by public officials, and praise for individuals acting in public interest. Note that neutral factual statements (e.g., “Netanyahu’s trial begins tomorrow in Jerusalem”) and descriptive mentions of corruption without a clear perspective will not be classified under this identity.
Leftist. Support for left-wing parties or policies, including dovish security policies, willingness to make territorial compromise, opposition to settlement construction, criticism of right-wing actors or policies, when tied to a clear ideological stance.
Note that:
  1. 1.
    Identification with left-wing parties or groups is labeled as Leftist.
  2. 2.
    This identity focuses on hawkish-dovish positions, not economic or social issues (which are coded separately, e.g., Liberal or Socially-oriented).
  3. 3.
    References to “peace” alone do not constitute a leftist identity without additional ideological context.
  4. 4.
    Centrist parties opposing a side are labeled according to the target of opposition (e.g., “Blue and White opposes Likud” = Leftist; “Blue and White opposes Labor” = Rightist).
  5. 5.
    Positions regarding state-religion relations are not considered part of the left/right ideological identities (e.g., referring to religious coercion would be considered liberal, not Leftist).
Rightist. Support for right-wing parties or policies, including hawkish security positions, Greater Israel ideology, criticism of left-wing actors or policies, when tied to a clear ideological stance.
Note that:
  1. 1.
    Identification with right-wing parties or groups is labeled as Rightist.
  2. 2.
    This identity focuses on hawkish-dovish positions, not economic or social issues (which are coded separately, e.g., Conservative or Capitalist).
  3. 3.
    References to Israel as a “Jewish and democratic state” do not alone indicate Rightist identity.
  4. 4.
    Centrist parties opposing a side are labeled according to the target of opposition (e.g., “Blue and White opposes Labor” = Rightist; “Blue and White opposes Likud” = Leftist;).
  5. 5.
    Positions regarding state-religion relations are not part of ideological identities (e.g., referring to greater congruence between state and religious laws would be considered Conservative not Rightist).
Palestinians and Arab Citizens of Israel. Statements regarding policy issues, worldviews, and ideologies that represent or reflect Palestinians and Arab-Israelis, including references to Palestinian and Arab-Israeli culture, statements and interviews by Palestinian and Arab-Israeli public leaders related to Palestinians and Arab-Israelis, as well as policy matters concerning these groups (not just security issues).
Note that:
  1. 1.
    References that do not reflect a Palestinian perspective (e.g., discussions on settlements or military actions against Palestinian organizations), will not be labeled under this identity.
  2. 2.
    The speaker does not have to be Palestinian but must present events or impacts from a Palestinian perspective.
  3. 3.
    Positive references for public figures representing these communities will be included under this identity.
  4. 4.
    This identity is distinct from the Leftist identity. References can be to both or either, depending on context.
Security-Oriented (Bitchonist). References to national security issues, military strength, security capabilities, threats to internal or external safety, defense agencies (IDF, Shin Bet, Mossad, etc.), borders protection, the state’s ability to protect its citizens and security challenges.
Note that:
  1. 1.
    This identity includes references that relate to the notion of protecting all citizens (e.g., Jewish, Arab-Israeli) from violence or terror.
  2. 2.
    This identity excludes general or symbolic mentions of soldiers unrelated to defense (e.g., prayers, greetings, daily life).
  3. 3.
    This identity includes critiques of external actors (e.g., referring to the Palestinian Authority as a terrorists funding organization) when framed in terms of national security.
Ultra-orthodox (Haredi). References to Jewish ultra-orthodox (Haredi) lifestyle, including the Haredian education system, gender segregation in the public sphere, exemption from military service for yeshiva students, charitable activities within the Haredi community, the Haredi religious (Torah) world and tradition preservation, Haredian parties (Shas, United Torah Judaism, Agudat Yisrael), and Haredi rabbinic leadership.
Note that:
  1. 1.
    The mere mentioning of Rabbis is not enough. Statements must carry a positive tone.
  2. 2.
    General religious expressions (e.g., quoting a verse, mentioning Jewish holidays) will not be labeled under this identity unless clearly framed within Haredian context or authority.
  3. 3.
    References to the integration of Haredim into broader Israeli society, or criticism of the Haredi community, will not be labeled under this identity.
Zionist. Affirmation of Zionist symbols and values, such as Jewish immigration to Israel (Aliyah), connections between Israel and the Jewish diaspora, national pride, IDF enlistment, national unity, symbolic expressions such as the positive references to the Israeli flag, national anthem and collective sacrifice (e.g., references to Memorial Day, families of fallen soldiers). Also includes references relating to Israel’s struggle against antisemitism and BDS.
Note that:
  1. 1.
    This definition draws on the Declaration of Independence, which defines Israel as the homeland of the Jewish people.
  2. 2.
    This identity includes expressions of dedication to Israel’s future, strength and prosperity (“to serve the country,” “for the future of the country”).
Appendix D Fine-tuning setup
All model fine-tuning is performed on an 80GB A100 Nvidia GPU, using huggingface transformers.
For decoder fine-tuning, the separator "### Answer:" is used to separate the input sentences from the output. The labels of the input and separator are loss-masked.
All experiments utilize AdamW optimizer with a linear scheduler. Default values of hyperparameters are used everywhere except for learning rate (for encoders and for decoder LLMs) and loss type (for encoder models).
Decoder fine-tuning uses QLORA with 4bit quantization. LORA settings are rank of 256, and alpha value of 512. LORA layers are attached to all linear levels in the decoder models.
Each hyper-parameter configuration was trained once.
Appendix E Generalization to Parliamentary Speech
Table 5 shows the full results of the 500 items sampled from parliamentary Knesset speeches for cross-genre generalization test. The predictions are produced by the best performing training checkpoint (DictaLM2.0 fine-tune).
Class Precision Recall F1-score Support
Conservative 0.33 0.82 0.47 17
Rightist 0.66 0.74 0.70 47
Democrat 0.76 0.69 0.72 74
Honest 0.72 0.81 0.76 47
Capitalist 0.68 0.87 0.76 31
Ultra-orthodox 0.66 0.84 0.74 32
Socially-oriented 0.73 0.72 0.73 72
Liberal 0.60 0.60 0.60 63
Leftist 0.66 0.77 0.71 53
Security-oriented 0.78 0.75 0.77 48
Palestinian 0.91 0.86 0.88 69
Zionist 0.76 0.81 0.78 42
Micro avg 0.70 0.76 0.73 595
Macro avg 0.69 0.77 0.72 595
Weighted avg 0.72 0.76 0.73 595
Samples avg 0.62 0.64 0.62 595
Table 5: Knesset sample results – DictaLM2.0 fine-tuned checkpoint
Appendix F Data Release
The annotated data is released under cc-by-4.0 license. The data is publicly available on github at https://github.com/guymorlan/hebid/.
Appendix G Annotation
All data has been annotated by two authors of the paper. The authors have not received direct compensation. The two annotators are Hebrew speaking women living in Israel.
Appendix H Model Sizes
Model Parameters
mBERT (bert-base-multilingual-cased) 110 M
AlephBERT 110 M
HeRo 125 M
DictaBERT–base 184 M
DictaBERT–large 340 M
Gemma–2B 2 B
Gemma–9B 9 B
Qwen3–8B 8.2 B
DictaLM 2.0 7 B
Table 6: Model sizes (number of parameters) for all models used in this paper.
Appendix I Preprocessing
Sentence segmentation was performed using the Stanza package.
Appendix J External Validity
CHES variable (2021) Correlated with Pearson r p-value
Overall Ideology (Left/Right) Rightist - Leftist 0.941 << 0.001
Rightist 0.880 0.002
Leftist -0.815 0.007
Social Values (Libertarian/Traditional) Conservative - Liberal 0.817 0.007
Conservative 0.431 0.247
Liberal -0.838 0.005
Civil Liberties vs. Law & Order Conservative - Liberal 0.905 << 0.001
Conservative 0.850 0.004
Liberal -0.845 0.004
Economic Stance (Left/Right) Capitalist - Socially-oriented 0.423 0.256
Capitalist 0.869 0.002
Socially-oriented 0.012 0.976
State’s Identity (Democratic vs. Jewish) Ultra-orthodox - Democrat 0.708 0.033
Ultra-orthodox 0.649 0.059
Democrat -0.410 0.274
Salience of Reducing Corruption Honest 0.664 0.051
Stance on Israeli-Palestinian Conflict Zionist 0.865 0.003
Palestinian -0.602 0.086
Position on a Palestinian State Zionist 0.862 0.003
Palestinian -0.549 0.126
Position on Engagement with Arab World Security-oriented 0.710 0.032
Table 7: Pearson correlation between CHES variables (2021) and identity discourse shares.
Appendix K Additional Results
Figure 7: Identity Correlation Matrix - Facebook
Figure 8: Identity Correlation Matrix - Knesset
Figure 9: Identity Correlation Matrix - Survey
Figure 10: Identity Ranking
Figure 11: Per-identity time-trends on Facebook (biweekly mean) and survey

### v3

@@
 1 Introduction
-Social identities—such as political ideology, religious affiliation, or demographic group membership—are powerful drivers of political behavior and public discourse (Tajfel and Turner, 1979).
-Yet existing NLP resources for identity detection remain almost entirely English‐focused, single‐label, and rely on coarse-grained categories (e.g., party or ethnicity).
-In this work, we introduce HebID, the first publicly released Hebrew dataset for fine‐grained, multilabel social identity detection in political text, grounded in both domain expertise and large‐scale survey evidence.
+Social identities—such as political ideology, religious affiliation, or demographic group membership—are powerful drivers of political behavior and public discourse (Tajfel et al., 2001).
+Yet existing NLP resources for identity detection remain almost entirely English-focused, single-label, and rely on coarse-grained categories (e.g., party or ethnicity).
+In this work, we introduce HebID, the first publicly released Hebrew dataset for fine-grained, multilabel social identity detection in political text, grounded in both domain expertise and large-scale survey evidence.
 We frame the task as a sentence-level classification problem: the goal is to determine which, if any, of twelve social identities are being positively expressed or invoked within a given sentence.
@@
 We utilized 12 waves of a national survey (N = 1,769) each of which included questions designed to identify the social identities most salient to Israeli citizens (e.g., Rightist, Ultra-Orthodox, Socially-oriented).
-These survey‐derived categories were then used to guide the annotation of 5,536 sentences sampled from Facebook posts by Israeli politicians (December 2018–April 2021), ensuring that our labels reflect real‐world identity salience.
-We benchmark a suite of modern architectures on the Facebook data—(i) multilabel encoder models; (ii) single‐label encoder models; and (iii) 2B–9B‐parameter decoder LLMs, finding that Hebrew‐tuned decoder models (specifically, DictaLM2.0) achieve the highest macro‐F1 (0.743), outperforming encoder‐only baselines by over six points.
-We then assess cross‐genre generalization by applying our best model to 500 parliamentary speech excerpts, achieving a comparable macro‐F1 of 0.72.
+These survey-derived categories were then used to guide the annotation of 5,536 sentences sampled from Facebook posts by Israeli politicians (December 2018–April 2021), ensuring that our labels reflect real-world identity salience.
+We benchmark a suite of modern architectures on the Facebook data—(i) multilabel encoder models; (ii) single-label encoder models; and (iii) 2B–9B-parameter decoder LLMs, finding that Hebrew-tuned decoder models (specifically, DictaLM2.0) achieve the highest macro-F1 (0.743), outperforming encoder-only baselines by over six points.
+We then assess cross-genre generalization by applying our best model to 500 parliamentary speech excerpts, achieving a comparable macro-F1 of 0.72.
 Finally, we link three complementary sources in our analysis of how identities behave: politicians’ Facebook posts, parliamentary speeches from the Israeli parliament (Knesset), and survey responses that reflect public identification.
-We (1) compare identity prevalence and correlations between social media, parliamentary speech, and the public; (2) document how identity discourse surges around election cycles; (3) uncover coherent “bundles” of co‐occurring identities; and (4) quantify gender‐related variation in identity expression.
+We (1) compare identity prevalence and correlations between social media, parliamentary speech, and the public; (2) document how identity discourse surges around election cycles; (3) uncover coherent “bundles” of co-occurring identities; and (4) quantify gender-related variation in identity expression.
 Our contributions are:
@@
 Group reference detection.
-Early work extended named‐entity recognition to capture social groups as entities.
+Early work extended named-entity recognition to capture social groups as entities.
 Zanotto et al.
-(2024) introduce GRIT, annotating Italian news and parliamentary text for spans referring to demographic, national, or partisan groups, and fine‐tune BERT to identify them.
-Licht and Sczepanski (2024) apply a similar span‐labeling approach to British parliamentary debates, quantifying how often politicians mention particular social groups.
+(2024) introduce GRIT, annotating Italian news and parliamentary text for spans referring to demographic, national, or partisan groups, and fine-tune BERT to identify them.
+Licht and Sczepanski (2024) apply a similar span-labeling approach to British parliamentary debates, quantifying how often politicians mention particular social groups.
 These studies demonstrate the feasibility of automatic group mention detection, but remain single-label and limited to explicit group mentions.
@@
 Them corpus (Huguet Cabot et al., 2021) annotates Reddit comments for target group, stance (supportive vs.
-hostile), and emotion, using a multi‐task RoBERTa model.
+hostile), and emotion, using a multi-task RoBERTa model.
 Card et al.
@@
 The Media Frames Corpus (Card et al., 2015) provides frame annotations (including cultural identity frames) for news articles.
-These resources focus on single‐label stance or framing, and predominantly English texts.
+These resources focus on single-label stance or framing, and predominantly English texts.
 Ideological position inference.
@@
 (2014) develop neural networks to predict left/right alignment of speeches.
-More recent prompt‐based methods with LLMs (Mishra et al., 2023) achieve zero‐shot ideological scoring.
+More recent prompt based methods with LLMs (Wang et al., 2022) achieve zero-shot ideological scoring (Le Mens and Gallego, 2025).
 While these approaches infer broad ideological leanings, they do not detect explicit identity mentions or support multilabel identity categories.
-In‐group/out‐group rhetoric.
-Discourse‐level signals such as pronoun clusivity and coded language reveal populist identity appeals.
+In-group/out-group rhetoric.
+Discourse-level signals such as pronoun clusivity and coded language reveal populist identity appeals.
 Rehbein and Ruppenhofer (2022) annotate inclusive vs.
-exclusive uses of “we” in German parliamentary debates and train classifiers to disambiguate referents, highlighting us‐versus‐them framing.
+exclusive uses of “we” in German parliamentary debates and train classifiers to disambiguate referents, highlighting us-versus-them framing.
 Mendelsohn et al.
 (2023) curate a dogwhistle lexicon and show that pretrained models struggle to detect covert slurs without knowledge grounding.
-These works emphasize the need for high‐precision, context‐aware annotation, but do not cover multilabel identity categories across many classes.
+These works emphasize the need for high-precision, context-aware annotation, but do not cover multilabel identity categories across many classes.
 Gaps and Our Contributions.
-Current identity‐focused corpora are typically single‐label, English‐centric, and restricted to broad political categories or explicitly stated group mentions.
-They rarely offer the fine‐grained, multilabel annotations needed to capture the complexity of real‐world political speech, and they do not cover non‐Latin scripts or languages such as Hebrew.
-We introduce the first publicly available Hebrew dataset for social identity detection, comprising 5,536 sentences from politicians’ Facebook posts annotated with twelve distinct identities—each grounded in survey‐measured salience and expert‐defined categories.
-By providing multilabel annotations across a rich set of ideologically, religiously, and socio‐economically relevant identities, our resource enables more nuanced analyses of how political actors deploy identity language than has previously been possible in any non‐English setting.
+Current identity-focused corpora are typically single-label, English-centric, and restricted to broad political categories or explicitly stated group mentions.
+They rarely offer the fine-grained, multilabel annotations needed to capture the complexity of real-world political speech, and they do not cover non-Latin scripts or languages such as Hebrew.
+We introduce the first publicly available Hebrew dataset for social identity detection, comprising 5,536 sentences from politicians’ Facebook posts annotated with twelve distinct identities—each grounded in survey-measured salience and expert-defined categories.
+By providing multilabel annotations across a rich set of ideologically, religiously, and socio-economically relevant identities, our resource enables more nuanced analyses of how political actors deploy identity language than has previously been possible in any non-English setting.
 3 Annotating Social Identities
@@
 Mean 96.1% 0.769
-Table 1: Inter‐Annotator Agreement (IAA) scores
+Table 1: Inter-Annotator Agreement (IAA) scores
 4.2 Descriptive statistics
@@
 We examine a set of encoder models and LLM decoders in the 2B–9B parameter range.
-For encoders, we use the multilingual mBERT (Devlin et al., 2019), and Hebrew-targeted encoders AlephBERT (Seker et al., 2022), HeRO (Hebron et al., 2023) and the base and large variants of DictaBERT (Shmidman et al., 2023).
+For encoders, we use the multilingual mBERT (Devlin et al., 2019), and Hebrew-targeted encoders AlephBERT (Seker et al., 2022), HeRO (Shalumov and Haskey, 2023) and the base and large variants of DictaBERT (Shmidman et al., 2023).
 All encoder models are trained for up to 10 epochs with three learning rates (1e-5, 3e-5, 5e-5), and three types of loss (default, positive weight and focal loss), choosing the best performing checkpoint on the validation set.
@@
 Macro-averaged metrics Per-label F1 scores
-Model P R F1
-Security-
-oriented
-Capitalist
-Conserva
--tive
-Democrat
-Ultra-
-Orthodox
-Socially-
-oriented
-Liberal
-Palestinian
-Leftist
-Rightist
-Honest
-Zionist
-Decoder‐only
+Model P R F1 Security- oriented Capitalist Conserva -tive Democrat Ultra- Orthodox Socially- oriented Liberal Palestinian Leftist Rightist Honest Zionist
+Decoder-only
 DictaLM2.0 0.740 0.751 0.743 0.705 0.805 0.675 0.754 0.653 0.723 0.724 0.852 0.750 0.765 0.816 0.700
-Gemma-2‐9B 0.717 0.698 0.705 0.645 0.759 0.684 0.753 0.591 0.662 0.749 0.758 0.671 0.737 0.804 0.650
-Gemma-2‐2B 0.620 0.631 0.624 0.560 0.723 0.575 0.757 0.385 0.634 0.673 0.716 0.609 0.615 0.680 0.557
-Qwen‐8B 0.665 0.463 0.542 0.605 0.508 0.308 0.700 0.410 0.541 0.568 0.625 0.516 0.515 0.650 0.554
+Gemma-2-9B 0.717 0.698 0.705 0.645 0.759 0.684 0.753 0.591 0.662 0.749 0.758 0.671 0.737 0.804 0.650
+Gemma-2-2B 0.620 0.631 0.624 0.560 0.723 0.575 0.757 0.385 0.634 0.673 0.716 0.609 0.615 0.680 0.557
+Qwen-8B 0.665 0.463 0.542 0.605 0.508 0.308 0.700 0.410 0.541 0.568 0.625 0.516 0.515 0.650 0.554
 Multilabel encoders
@@
 mBERT 0.573 0.553 0.552 0.483 0.467 0.527 0.708 0.356 0.568 0.635 0.632 0.531 0.502 0.653 0.566
-Single‐label encoders
+Single-label encoders
 DictaBERTLarge 0.643 0.689 0.659 0.672 0.659 0.529 0.731 0.508 0.708 0.694 0.783 0.615 0.636 0.721 0.652
@@
 LogisticRegression 0.469 0.268 0.334 0.350 0.000 0.280 0.500 0.280 0.410 0.340 0.470 0.380 0.350 0.360 0.290
-Table 4: Macro‐averaged precision (P), recall (R), F1, and per‐label F1 for all models.
+Table 4: Macro-averaged precision (P), recall (R), F1, and per-label F1 for all models.
 Table 4 reports test set results for trained models.
@@
 This work was supported by the Israel Science Foundation (Grant no.
-2315/18).
+2501/22).
 We thank Vered Porzycki for assistance in conceptualizing the definitions of the annotated identities; Vered Porzycki, Gal Ron, Avishai Green, Dror Markus and Noa Amir for invaluable assistance in a preliminary pilot study; and Effi Levi for his very useful and insightful advice.